Suppose the MPC is .9 and planned investment increases by $100 billion. After the multiplier process has taken effect, the Aggregate Demand curve will have shifted to the right by $_______________ billion.
Answer #1

Answer

Change in AD =change in investment * multiplier

multiplier =1/(1-MPC)

=1/(1-0.9)

=10

Change in AD =100*10

=$1000 billion

the GDP shifts right by $1000 billion